OpenStudy (anonymous):

what is 1+2+3+...........+999999998+999999999+1000000000= ?

OpenStudy (anonymous):

sum(1 to n) = (1/2)(n)(n+1) e.g., sum (1 to 10) = (1/2)(10)(11)= 55

OpenStudy (anonymous):

cool what subject did you learn that in?

OpenStudy (anonymous):

I discovered it myself in high school. Found later was first done by Gauss centuries ago. Thanks for asking.

OpenStudy (anonymous):

nice thanks for that